\begin{thebibliography}{99}

\bibitem{QC}  K. Abdukhalikov, T. Bag and D. Panario, One-generator quasi-cyclic codes and their dual codes, Discrete Math. 346(6),  113369, (2023). 

\bibitem{0} S. A. Aly, A. Klappenecker and P. K. Sarvepalli, On quantum and classical BCH codes, IEEE Trans. Inf. Theory 53(3),  1183--1188, (2007).



\bibitem{A} A. Ashikhmin and E. Knill, Nonbinary quantum stabilizer codes, IEEE Trans. Inf. Theory 47,  3065--3072, (2001).

\bibitem{Nuh1} N. Aydin, I. Siap and D. K. Ray-Chaudhuri, The structure of 1-generator quasi-twisted codes and new linear codes, Des. Codes Cryptogr. 24,  313--326, (2001).

\bibitem{Mag} W. Bosma, J.  Cannon, C. Fieker and A. Steel (eds.),  Handbook of magma functions, Edition 2.19,  5488 pages, (2013).

\bibitem{iitr}
S. Benjwal, M. Bhaintwal,  On the duals of quasi-cyclic codes and their application to quantum codes, Quantum Inf. Process 23, 113, (2024).


\bibitem{cal1996}
A. R. Calderbank and P. W. Shor, Good quantum error-correcting codes exist, Phys. Rev. A 54(2), 1098--1105, (1996).

\bibitem{CRSS98} A. R. Calderbank, E. M. Rains, P. M. Shor and N. J. A. Sloane,   Quantum error-correction via codes over $\text{GF}(4)$, IEEE Trans. Inf. Theory  44,  1369--1387, (1998).





\bibitem{Con} J. Conan and G. Seguin,   Structural properties and enumeration of quasi cyclic codes,  Appl. Algebra Eng. Commun. Comput. 4,  25--39, (1993).


\bibitem{Gal} C. Galindo, F. Hernando and R. Matsumoto, Quasi-cyclic constructions of quantum codes, Finite Fields Appl.  52,  261--280, (2018).

\bibitem{gra}
M. Grassl, Bounds on the minimum distance of linear codes and quantum codes, Online available at http://www.codetables.de,  Accessed on 2025-05-17.

\bibitem{Q4}
C. Guan, R. Li and L. Lu,  New binary quantum codes constructed from quasi-cyclic codes, Int. J. Theor. Phys 61, 172, (2022).


\bibitem{Q1}C. Guan, R. Li, J. Lv and Z. Ma, Symplectic self-orthogonal quasi-cyclic codes, ArXiv:2212.14225.





\bibitem{Kas} 
T. Kasami, A Gilbert-Varshamov bound for quasi-cyclic codes of rate $\frac{1}{2}$, IEEE Trans. Inf. Theory 20,  679, (2018).

\bibitem{Ket} A. Ketkar, A. Klappenecker, S. Kumar, and P. K. Sarvepalli, Nonbinary stabilizer codes over finite fields, IEEE Trans. Inf. Theory 52,  4892–-4914, (2006).



\bibitem{La} K. Lally and P. Fitzpatrick, Algebraic structure of quasicyclic codes, Discrete Appl. Math. 11,  157--175, (2001).

\bibitem{L1} S. Ling and P. Sol\'e,  On the algebraic structure of quasi-cyclic codes I: finite fields, IEEE Trans. Inf. Theory  47(7),  2751--2760, (2001).


\bibitem{L5} S. Ling and P. Sol\'e, Good self-dual quasi-cyclic codes exist, IEEE Trans. Inf. Theory 49(4), 1052--1053, (2003). 



\bibitem{L2} S. Ling and P. Sol\'e,  On the algebraic structure of quasi-cyclic codes II: chain rings, Des. Codes Cryptogr. 30,  113--130, (2003).


\bibitem{L3} S. Ling and P. Sol\'e,  On the algebraic structure of quasi-cyclic codes III: generator theory,  IEEE Trans. Inf. Theory  51(5), 2692--2700, (2005).



\bibitem{L4} S. Ling and P. Sol\'e,  On the algebraic structure of quasi-cyclic codes IV: repeated roots, Des. Codes Cryptogr. 38,  337--361, (2006).



\bibitem{Q3} J. Lv, R Li and J. Wang,  New binary quantum codes derived from one-generator quasi-cyclic codes, IEEE Access 7,  85782--85785, (2019).

\bibitem{Q2}
J. Lv, R. Li and J. Wang, An explicit construction of quantum stabilizer codes from quasi-cyclic codes, IEEE Commun. Lett.  24(5), 1067--1071, (2020).


\bibitem{Se}  G. E. Seguin, A class of 1-generator quasi-cyclic codes, IEEE Trans. Inf. Theory 50,  1745--1753, (2004).



\bibitem{St} A. M. Steane,  Simple quantum error-correcting codes, Phys. Rev. A 54, 4741--4751, (1996).

\bibitem{S95} P. W. Shor,  Scheme for reducing decoherence in quantum memory, Phys. Rev. A  52,  2493--2496, (1995).



\bibitem{xu2022}
H. Xu and W. Du, On some binary symplectic self-orthogonal codes, Appl. Algebra Eng. Commun. Comput. 33, 321--337, (2022).


\end{thebibliography}

\documentclass[11pt, a4paper]{article}

\usepackage{amsmath, amsfonts, amssymb, amsthm}
\usepackage{graphicx}
\usepackage{epsfig}
\usepackage{float}
\usepackage{color}
\usepackage[top=1in, bottom=1in, left=1in, right=1in]{geometry}
\usepackage{url} 
\usepackage{comment}
\usepackage[linesnumbered,ruled,vlined]{algorithm2e}

\usepackage[pagewise,mathlines]{lineno} 

\usepackage{amsmath,amssymb,amsfonts}

\SetCommentSty{mycommfont}

\DeclareMathOperator{\lcm}{lcm}

\DeclareMathOperator{\rank}{rank}
\numberwithin{figure}{section}
\newtheorem{theorem}{Theorem}[section]
\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{corollary}[theorem]{Corollary}
\newtheorem{proposition}[theorem]{Proposition}
\newtheorem{example}[theorem]{Example}
\newtheorem{remark}[theorem]{Remark}
\newtheorem{definition}[theorem]{Definition}

\begin{document}
\title{Some New Non-binary Quantum Codes from One-generator  Quasi-cyclic Codes}
\author{Tushar Bag$^{1,2}$, Hai Q. Dinh$^3$ , Daniel Panario$^4$}
\date{
\small{
1. Department of Mathematics, SRM University-AP, Amaravati, Andhra Pradesh 522240, India\\
2. Inria, ENS de Lyon, Lyon 1, LIP, 69342, Lyon cedex 07, France\\
3. Kent State University, Warren, OH 44483, USA\\
4. School of Mathematics and Statistics, Carleton University, Ottawa, Ontario K1S 5B6, Canada.
}
\today}

\begingroup
\renewcommand\thefootnote{}
\footnotetext{
Email: tusharbag2011@gmail.com (T. Bag) [corresponding author], 
hdinh@kent.edu (H. Q. Dinh), 
daniel@math.carleton.ca (D. Panario)
}
\endgroup

\maketitle

\begin{abstract}
This article studies one-generator and two-generator quasi-cyclic codes over finite fields. We present necessary and sufficient conditions for the symplectic self-orthogonality of one-generator quasi-cyclic codes in two equivalent forms: a matrix form and a polynomial form. We then extend these conditions to two-generator quasi-cyclic codes, providing analogous characterizations for both the symplectic self-orthogonality and the symplectic dual-containing property. Using these conditions, we construct new quantum error-correcting codes whose parameters improve upon the current best-known records.
\end{abstract}
\text{\bf Keywords:} \small{Quasi-cyclic codes (QCs),  quantum error-correcting codes (QECCs).}\\
\text{\bf Mathematics Subject Classifications(2010):} \small{ 94B05, 94B15, 94B60.}\\

\section{Introduction}
Quasi-cyclic (QC) codes are a prominent class of linear error-correcting codes. They possess a well-developed algebraic structure that generalizes  cyclic codes. This generalization allows greater flexibility in code design and enables the construction of asymptotically good codes  approaching the modified Gilbert--Varshamov bound \cite{Kas, L5}. The study of QC codes has yielded numerous record-breaking linear codes, particularly over small finite fields. Several key contributions have been made to the understanding of QC codes. Conan et al.\ \cite{Con} studied the structural properties of quasi-cyclic codes, providing both an enumeration of these codes and a characterization of their duals. Seguin \cite{Se} examined the properties of a specific class of one-generator QC codes. Ling and Sol\'e investigated the algebraic structure of QC codes in a series of articles \cite{L1, L2, L3, L4}. Lally et al.\ \cite{La} explored the structure and duals of arbitrary QC codes, with a particular focus on self-dual QC codes of index of 2. Aydin et al. \cite{Nuh1} investigated the structure of $1$-generator quasi-twisted codes and constructed new linear codes with improved parameters.

\vskip 3pt

Quantum error-correcting codes (QECCs) are essential for protecting quantum information from decoherence and quantum noise, playing a significant role in both quantum computing and communication. Quantum computers are expected to solve certain classes of problems significantly faster than classical computers. However, mitigating errors caused by decoherence and noise remains a critical challenge, and QECCs have emerged as a powerful tool for safeguarding quantum information in both communication and computation. The concept of quantum error correction was first proposed independently in \cite{cal1996, St, S95}.  The Calderbank--Shor--Steane (CSS) construction \cite{CRSS98} has served as the foundation for a large portion of subsequent research on QECCs.  Techniques for constructing non-binary quantum codes are explored in \cite{0, A, Ket}. 

\vskip 3pt

Symplectic self-orthogonal quasi-cyclic (QC) codes have proven to be an excellent family not only for constructing new linear codes but also for constructing numerous new binary QEECs. The study of QECCs construction from QC codes gained momentum following the work of Galindo et al.\ \cite{Gal}, where the authors studied a specific class of $2$-generator QC codes using Euclidean, Hermitian, and symplectic inner products. 
Subsequently, Lv et al.\ \cite{Q3, Q2} and Guan et al.\ \cite{Q4, Q1} constructed several record-breaking binary quantum codes by exploiting the symplectic structure of QC codes. Explicit dual generators of QC codes have been studied in \cite{QC, iitr}.

\vskip 3pt
The motivation for this work is to derive necessary and sufficient conditions for the symplectic self-orthogonality and the symplectic dual-containing property in a unified and simplified form that applies to general quasi-cyclic codes. While specific forms have been studied previously in the literature, we aim to subsume these as special cases, provide a simpler and more general formulation, and demonstrate that our approach yields new codes with record-breaking parameters.
\vskip 3pt

This paper is organized as follows. In Section 2, we present the basics of linear codes and quasi-cyclic codes. In Section 3, we study one-generator quasi-cyclic codes and present the symplectic self-orthogonality condition over finite fields. Section 4 focuses on two-generator quasi-cyclic codes, showing both the symplectic self-orthogonality condition and the symplectic dual-containing condition for the general form of two-generator quasi-cyclic codes over finite fields. In Section 5, we construct new quantum codes with record-breaking parameters based on our study. Finally in Section 6, we conclude our work giving further research problems.

\section{Preliminaries}

Let \(\mathbb{F}_q\) be the finite field with \(q = p^r\) elements, where \(p\) is a prime number and \(r\) is a positive integer. A linear code $C$ of length $2n$ over $\mathbb{F}_q$ is a subspace of the vector space $\mathbb{F}_q^{2n}$. The elements of \(C\) are called codewords. Suppose \(\mathbf{a} = (a_0, a_1, \dots, a_{2n-1})\) and \(\mathbf{b} = (b_0, b_1, \dots, b_{2n-1})\) are codewords of \(C\). 
The minimum Hamming weight of $C$ is defined as
\[w_H(C) = \min \{ w_H(\mathbf{a}) \mid \mathbf{a} \in C,\, \mathbf{a} \ne \mathbf{0} \},\]
where $w_H(\mathbf{a})$ denotes the number of nonzero components of $\mathbf{a}$.

The (minimum) Hamming distance of \(C\) is $ d_H(C) = \min \{ d_H(\mathbf{a}, \mathbf{b}) \mid \mathbf{a},\mathbf{b}\in C, \mathbf{a} \ne \mathbf{b} \},$
where \( d_H(\mathbf{a}, \mathbf{b}) = |\{ i \mid a_i \ne b_i \}|\).
\vskip 5pt
The symplectic inner product of \(\mathbf{u} = (\mathbf{u}_1, \mathbf{u}_2) \in \mathbb{F}_q^{2n}\) and \(\mathbf{v} = (\mathbf{v}_1, \mathbf{v}_2) \in \mathbb{F}_q^{2n}\) is defined as
\[ \langle \mathbf{u}, \mathbf{v} \rangle_s = \langle \mathbf{u}_1, \mathbf{v}_2 \rangle_e - \langle \mathbf{v}_1, \mathbf{u}_2 \rangle_e, \]
where \(\mathbf{u}_1, \mathbf{u}_2, \mathbf{v}_1, \mathbf{v}_2 \in \mathbb{F}_q^n\) and \(\langle \cdot, \cdot \rangle_e\) is the standard Euclidean inner product in \(\mathbb{F}_q^n\). We observe that this inner product can also be written as
\[ \langle \mathbf{u}, \mathbf{v} \rangle_s = \mathbf{u} \Omega \mathbf{v}^t, \]
where
\[ \Omega = \begin{pmatrix} O_n & I_n \\ -I_n & O_n \end{pmatrix}, \]
where \( I_n \) denotes the \( n \times n \) identity matrix, and \( O_n \) denotes the \( n \times n \) zero matrix.
\vskip 5pt
The symplectic dual code \(C^{\perp_s}\) of \(C\) is defined as $ C^{\perp_s} = \{ \mathbf{u} \in \mathbb{F}_q^{2n} \mid \langle \mathbf{u}, \mathbf{v} \rangle_s = 0, \text{ for all } \mathbf{v} \in C \}. $
A linear code \(C\) is called symplectic self-orthogonal if \(C \subseteq C^{\perp_s}\), and symplectic dual-containing if \(C^{\perp_s} \subseteq C\).
Let \(\mathbf{c} = (\mathbf{x}, \mathbf{y}) \in \mathbb{F}_q^{2n}\), where \(\mathbf{x}, \mathbf{y} \in \mathbb{F}_q^n\). We define the (minimum) symplectic weight of \(C\) as $ w_S(C) = \min \{ w_S(\mathbf{c}) \mid \mathbf{c} \in C, \mathbf{c} \ne 0 \}, $
where \(w_S(\mathbf{c}) = w_S(\mathbf{x}, \mathbf{y}) = |\{ i \mid (x_i, y_i) \ne (0,0) \}|\).

\vskip 3pt

A linear code $C$ of length $n$ over $\mathbb{F}_q$ is  a \emph{cyclic code} if for every codeword $(c_0,c_1,\dots,c_{n-1})\in C$, the cyclic shift $(c_{n-1},c_0,c_1,\dots,c_{n-2})$ is also in $C$. Let $R=\mathbb{F}_q[x]/\langle x^n-1 \rangle$. We can identify a codeword $(c_0,c_1,\dots,c_{n-1})\in C$ by a polynomial $c(x)=c_0+c_1x+\cdots+c_{n-1}x^{n-1}\in R$. It is easy to show that   $C$ is a cyclic code of length $n$ over $\mathbb F_q$ if it forms an ideal of $R$.

\begin{definition}
Suppose $C$ is a linear code of length $ln$ over $\mathbb F_q$. Any codeword ${\mathfrak c}=(c_{0,0},c_{0,1},\dots,\linebreak  c_{0,l-1},c_{1,0},\dots, c_{1,l-1}, \dots, c_{n-1,0},\dots,c_{n-1,l-1}) \in C$  can be written   as 

\[{\mathfrak c}=\begin{pmatrix}
c_{0,0} & c_{0,1} &\cdots &c_{0,l-1}\\
c_{1,0} & c_{1,1} &\cdots &c_{1,l-1}\\
\vdots & \vdots & \vdots &\vdots \\
c_{n-1,0} & c_{n-1,1} &\cdots &c_{n-1,l-1}\\
\end{pmatrix}.
\]
In this case,  $C$ is  a quasi-cyclic (QC) code of index $l$ if for any 
${\mathfrak c}\in C$, we get

\[\begin{pmatrix}
c_{n-l,0} & c_{n-l,1} &\cdots &c_{n-l,l-1}\\
c_{0,0} & c_{0,1} &\cdots &c_{0,l-1}\\
\vdots & \vdots & \vdots &\vdots \\
c_{n-l-1,0} & c_{n-l-1,1} &\cdots &c_{n-l-1,l-1}\\
\end{pmatrix}\in C.\]
\end{definition}

\section{One-generator QC codes}
A QC code of length $2n$ and index $2$ can be written as $C = (C_1, C_2)$, where $C_j$ is a cyclic code of length $n$ for $j=1,2$.
Suppose \(C_j\) is generated by a polynomial \(c_j(x)\) such that \(c_j(x) \mid x^n - 1\) for \(j = 1, 2\). Then, a one-generator QC code with index \(2\) can be interpreted as a 2-tuple of polynomials \((c_1(x), c_2(x))\).
\vskip 3pt
Any one-generator QC code of length \(2n\) and index \(2\) can be written as
\[
C = \big\{r(x)\big(c_1(x), c_2(x)\big) \mid r(x) \in R\big\} = \big\{\big(r(x)c_1(x), r(x)c_2(x)\big) \mid r(x) \in R\big\}.
\]

\begin{theorem}\label{gen}
Let $C$ be a one-generator QC code of length $2n$ and index $2$ over $\mathbb F_q$. Then a generator of $C$ is of the form $(r_1(x)g_1(x), r_2(x)g_2(x))$, 
where ${g_i}(x)h_i(x)= x^n -1$ and $\gcd\big(r_i(x),h_i(x)\big)=1$ for  $i = 1, 2$.
\end{theorem}

\begin{proof}
    Let \( C \) be a one-generator QC code of length \( 2n\) and index \( 2 \) over \( \mathbb{F}_q \), generated by \((c_1(x), c_2(x))\). Any element of \( C \) can be written in the form \(\big(s(x)c_1(x), s(x)c_2(x)\big)\) for some polynomial \( s(x) \in R \).
\vskip 3pt
For \( i = 1, 2 \), we define a map \(\Psi_i : C \longrightarrow R\) by
\[
\Psi_i\big(s(x)c_1(x), s(x)c_2(x)\big) = s(x)c_i(x).
\]

It is straightforward to verify that \(\Psi_i\) is a module homomorphism. 
Since \(\Psi_i(C)\) is the image of \( C \) under a module homomorphism, it forms an ideal of \( R \).
\vskip 3pt
Since every ideal of $R = \mathbb{F}_q[x] / \langle x^n - 1 \rangle$ corresponds to a cyclic code of length $n$ over $\mathbb{F}_q$, and $R$ is a principal ideal ring, the ideal $\Psi_i(C)$ is generated by a single polynomial $g_i(x)$ with $g_i(x) \mid x^n - 1$.
\vskip 3pt
Thus, the code \( C \) can be expressed as
\[
C = \big(r_1(x)g_1(x), r_2(x)g_2(x)\big) ,
\]
where \( g_i(x) h_i(x) = x^n - 1 \) and \(\gcd(r_i(x), h_i(x)) = 1\) for \( i = 1, 2 \). This completes the proof.
\end{proof}

\begin{definition}
Let $C$ be a one-generator QC code of length $2n$ and index $2$ over $\mathbb F_q$. Then the monic polynomial $h(x)$ of minimum degree such that
\[h(x)\big(r_1(x)g_1(x), r_2(x)g_2(x)\big)=(0,0)\]
\noindent is the parity-check polynomial of $C$.
\end{definition}

\begin{theorem}\label{gcd1}
Let $C=\big(r_1(x)g_1(x), r_2(x)g_2(x)\big)$ be a one-generator QC code  of length $2n$ and index $2$ over $\mathbb F_q$, where \( g_i(x) h_i(x) = x^n - 1 \) and \(\gcd(r_i(x), h_i(x)) = 1\) for \( i = 1, 2 \). Then $\dim(C)=\deg(h(x))$.

\end{theorem}

\begin{proof}
We define a module homomorphism \(\Phi: R \rightarrow C\) by 
\[\Phi(a(x)) = a(x)\big(r_1(x)g_1(x), r_2(x)g_2(x)\big).
\]

Let \( h(x) = \mathrm{lcm}(h_1(x), h_2(x)) \) be the parity-check polynomial of \(C\). We note that \(\text{ker}(\Phi) = (h(x))\).
Since  \(\Phi\) is surjective, we obtain \( R/(h(x)) \cong C \), which implies  $\dim(C) =\deg(h(x))$.
\end{proof}
\begin{remark}{\em
An important observation regarding  Theorem \ref{gcd1} is that, without the conditions \linebreak
\(\gcd(r_i(x), h_i(x)) = 1\) for \(i = 1, 2\), we cannot assert that \(\dim(C) = \deg(h(x))\). 
 We show this with the following example.\hfill$\square$
}\end{remark}

\begin{example}{\em
We consider \(R = \mathbb{F}_3[x]/\langle x^{10} - 1 \rangle\), and 
\[x^{10}-1 = (x+1)(x+2)(x^4 + x^3 + x^2 + x + 1)(x^4 + 2x^3 + x^2 + 2x + 1)\in \mathbb{F}_3[x].\]
 We take  \[g_1(x) =(x+2)(x^4 + 2x^3 + x^2 + 2x + 1),~~g_2(x) = (x^4 + x^3 + x^2 + x + 1)(x^4 + 2x^3 + x^2 + 2x + 1),\]
\[r_1(x) = 2x^4 + 2x^3 + 2x^2 + 2x + 2,~~\mbox{and}~~
r_2(x) = 2x^5 + 2x^4 + x^3 + x^2 + 2x.\]
\vskip 3pt
We have that $g_i(x) \mid x^n - 1$, and $r_i(x)\in R$, for \(i = 1, 2\). 
Also, as \( g_i(x) h_i(x) = x^n - 1 \) for \( i = 1, 2 \), we get
\[h_1(x) = x^5 + 2x^4 + 2x^3 + 2x^2 + 2x + 1,~~\mbox{and}~~
h_2(x) = x^2 + 2.\]
\vskip 3pt
From the above we get, 
\[\gcd(r_1(x),h_1(x)) = x^4 + x^3 + x^2 + x + 1,~~\mbox{and}~~
\gcd(r_2(x),h_2(x)) = 1. \]
Then $h(x)=\lcm(h_1(x),h_2(x))=x^6 + x^5 + 2x + 2$, implying $\deg(h(x))=6$. 
\vskip 4pt
On the other hand, using MAGMA computations, we find that the dimension of the QC code generated by \((r_1(x)g_1(x), r_2(x)g_2(x))\) is 2. 
\hfill$\square$
}\end{example}

\begin{corollary}\cite{Nuh1}
    Let $C=(r_1(x)g(x), r_2(x)g(x))$ be a one-generator QC code of length $2n$ and index $2$ over $\mathbb F_q$. Then $\dim(C)=n-\deg(g(x))$. Moreover,  $d'(C) \ge 2d_H$, where $d'$ is the minimum Hamming distance of $C$ and $d_H$ is the minimum Hamming distance of $C_i$ for $i=1,2$. \hfill$\square$
\end{corollary}

\begin{theorem}\label{mt}\cite{xu2022}
Let \( C \) be a linear code of length \( 2n \) over \( \mathbb{F}_q \) with generator matrix \( G \). Suppose \( G \) is an \( m \times 2n \) matrix. Then \( C \) is a symplectic self-orthogonal code if and only if \( G \Omega G^t = O_m \), where \( O_m \) is the \( m \times m \) zero matrix and \( G^t \) denotes the transpose of \( G \).
\end{theorem}

\begin{proof}
    Suppose \( C \) is a symplectic self-orthogonal, that is, \( C \subseteq C^{\perp_s} \). Then
    \begin{align*}
        C \subseteq C^{\perp_s}
       & \Longleftrightarrow \forall x \in C, \quad x \in C^{\perp_s}, \\
       & \Longleftrightarrow \forall x \in C, \quad G\Omega x^T = O_{m \times 1}, \\
       & \Longleftrightarrow \text{for all rows } r \text{ of } G, \quad G \Omega r^T = O_{m \times 1}, \\
       & \Longleftrightarrow G\Omega G^T = O_m.
    \end{align*}
\end{proof}

We recall a generator of a one-generator QC code of length \( 2n \) and index \( 2 \) over \( \mathbb{F}_q \) as \( C = (r_1(x)g_1(x), r_2(x)g_2(x)) \), and we denote  \( a(x) = r_1(x)g_1(x) \) and \( b(x) = r_2(x)g_2(x) \). We note that $a(x),  b(x) \in R$. Then a generator matrix of \( C \) can be expressed as \( G = (A \mid B) \), where \( A \) and \( B \) are \( n \times n \) circulant matrices generated by \( a(x) \) and \( b(x) \), respectively. Here, \( \mid \) denotes the horizontal concatenation of the two circulant matrices \( A \) and \( B \). Then we have the following result.

\begin{theorem}\label{t1}
Let \(C\) be a one-generator QC code of length $2n$ and index $2$ over \(\mathbb{F}_q\), whose generator matrix is \(G = (A \mid B)\). Then \(C \subseteq C^{\perp_s}\) if and only if \(AB^t = BA^t\), where \(A^t\) and \(B^t\) represent the transposes of \(A\) and \(B\), respectively.
\end{theorem}

\begin{proof}
    By Theorem \ref{mt}, \( C \subseteq C^{\perp_s} \) if and only if \( G \Omega G^t \) is a zero matrix. For this one-generator QC code, the generator matrix \( G \) is of the form \( G = (A \mid B) \). Thus, we have

\begin{align*}
G \Omega G^t &= \begin{pmatrix}
A & B 
\end{pmatrix}
\begin{pmatrix}
O_n & I_n \\
-I_n & O_n
\end{pmatrix}
\begin{pmatrix}
A & B 
\end{pmatrix}^t= \begin{pmatrix}
-B & A
\end{pmatrix}
\begin{pmatrix}
A^t  \\
B^t 
\end{pmatrix} =  - B A^t + A B^t.
\end{align*}
Therefore, $A B^t - B A^t = O_n$, implies $A B^t = B A^t $. Hence, $C\subseteq C^{\perp_s}$ if and only if $A B^t=B A^t$.
\end{proof}

We can also present Theorem \ref{t1} in terms of polynomials. To do so, we need to discuss the transpose of a polynomial, and its relation with the generator matrix described as follows.

\vskip 5pt

Let $t(x)=t_0 +t_1 x+t_2 x^2+\cdots+t_{n-2}x^{n-2}+t_{n-1}x^{n-1} \in \mathbb F_q[x]/\langle x^n-1 \rangle$. 
We define the transpose polynomial $\overline{t}(x)$ of $t(x)$ as 
\[\overline{t}(x) =x^n t(x^{-1}) =t_0 +t_{n-1} x+t_{n-2} x^2+\cdots+t_{2}x^{n-2}+t_{1}x^{n-1}. \]

\begin{lemma}
Let $b(x)=b_0+b_1x+\cdots+b_{n-1}x^{n-1}
\in \mathbb F_q[x]/\langle x^n-1\rangle,$
and let \(B\) be the circulant matrix generated by \(b(x)\).
Define
\[
\overline{b}(x)=x^n b(x^{-1})
=
b_0+b_{n-1}x+b_{n-2}x^2+\cdots+b_1x^{n-1}.
\]
Then the transpose matrix \(B^t\) is the circulant matrix generated by
\(\overline{b}(x)\).
\end{lemma}

\begin{proof}

The circulant matrix generated by $b(x)=b_0+b_1x+\cdots+b_{n-1}x^{n-1}$
is
\[
B=
\begin{pmatrix}
b_0 & b_1 & b_2 & \cdots & b_{n-1}\\
b_{n-1} & b_0 & b_1 & \cdots & b_{n-2}\\
b_{n-2} & b_{n-1} & b_0 & \cdots & b_{n-3}\\
\vdots & \vdots & \vdots & \ddots & \vdots\\
b_1 & b_2 & b_3 & \cdots & b_0
\end{pmatrix}.
\]
If rows and columns are indexed from \(0\) to \(n-1\), then $
B_{ij}=b_{j-i \pmod n}.$
Now we consider the transpose matrix \(B^t\). Its entries are
\[
(B^t)_{ij}=B_{ji}.
\]
Using the formula for the entries of \(B\), we get $
(B^t)_{ij}=b_{i-j \pmod n}.$ We write
\[
\overline{b}(x)
=
\sum_{k=0}^{n-1} c_k x^k,~\mbox{where}~ c_k=b_{-k \pmod n}.
\]
The circulant matrix generated by \(\overline{b}(x)\) has entries $c_{j-i}
=
b_{-(j-i)\pmod n}.$ Since
\[
-(j-i)\equiv i-j \pmod n,
\]
we get
\[
c_{j-i}=b_{i-j \pmod n}.
\]
Therefore, the \((i,j)\)-entry of the circulant matrix generated by
\(\overline{b}(x)\) is exactly $(B^t)_{ij}.$ Hence \(B^t\) is the circulant matrix generated by
\(\overline{b}(x)\).

\end{proof}

\begin{proposition} \label{prf0}
    Let \(A\) and \(B\) be circulant matrices generated respectively by $a(x), b(x)\in \mathbb F_q[x]/\langle x^n-1\rangle.$
Then the product  \(AB^t\) is the circulant matrix generated by $a(x)\overline{b}(x)\mod (x^n - 1).$
\end{proposition}

\begin{proof}
    Since \(A\) is generated by \(a(x)\), its entries satisfy $A_{ij}=a_{j-i \pmod n}.$
By the previous lemma, \(B^t\) is generated by \(\overline{b}(x)\), and
$(B^t)_{ij}=b_{i-j \pmod n}.$ Let $C=AB^t,$ the \((i,j)\)-entry of \(C\) is
\[
C_{ij}
=
\sum_{k=0}^{n-1}
A_{ik}(B^t)_{kj}.
\]
Substituting the formulas for the entries,
\[
C_{ij}
=
\sum_{k=0}^{n-1}
a_{k-i}\, b_{k-j},
\]
where all indices are taken modulo \(n\).

Let $m=k-i$, and set $r=j-i.$ Then
\[
k-j=(k-i)-(j-i)=m-r.
\]
Hence,
\[
C_{ij}
=
\sum_{m=0}^{n-1}
a_m b_{m-r}.
\]
Thus, \(C_{ij}\) depends only on \(r=j-i\), so \(C\) is circulant.

\vskip 5pt \noindent
Next, we compute the product $a(x)\overline{b}(x).$ Since $\overline{b}(x)
=\sum_{t=0}^{n-1} b_{-t}x^t,$ we have
\[
a(x)\overline{b}(x)
=
\left(\sum_{m=0}^{n-1} a_mx^m\right)
\left(\sum_{t=0}^{n-1} b_{-t}x^t\right).
\]
Multiplying we get 
\[
a(x)\overline{b}(x)
=
\sum_{m=0}^{n-1}\sum_{t=0}^{n-1}
a_m b_{-t} x^{m+t}.
\]
Modulo \(x^n-1\), the coefficient of \(x^r\) is
\[
\sum_{m+t~\equiv~ r \pmod n}
a_m b_{-t}.
\]
By setting $t=r-m,$ we get the coefficient of \(x^r\) as
\[
\sum_{m=0}^{n-1}a_m b_{m-r}.
\]
This is exactly the expression obtained for \(C_{ij}\) when
\(r=j-i\). Therefore the circulant matrix \(C=AB^t\) is generated by $a(x)\overline{b}(x)\pmod{x^n-1}.$
\end{proof}

\vskip 5pt
We present the following result on the symplectic self-orthogonality of a one-generator quasi-cyclic code in terms of the generator polynomials.

\begin{theorem}\label{prf}
Let \(C\) be a one-generator QC code of length $2n$ and index $2$ generated by \((a(x), b(x))\). Then \(C \subseteq C^{\perp_s}\) if and only if \(a(x)\overline{b}(x) - b(x)\overline{a}(x) \equiv 0 \mod (x^n - 1) \).
\end{theorem}

\begin{proof}
    By Theorem \ref{t1}, the condition \( C \subseteq C^{\perp_s} \) holds if and only if \( AB^t = BA^t \). We aim to express this condition using polynomials.
    By Proposition \ref{prf0}, the product matrix \( AB^t \) corresponds to the circulant matrix generated by \( a(x) \overline{b}(x) \mod (x^n - 1) \), and \( BA^t \) corresponds to the circulant matrix generated by \( b(x) \overline{a}(x) \mod (x^n - 1) \). Therefore, \(AB^t = BA^t\) if and only if \(a(x)\overline{b}(x) - b(x)\overline{a}(x) \equiv 0 \mod (x^n - 1)\).
\end{proof}

\vskip 3pt
Here, we present a detailed example explaining all the concepts discussed above.

\begin{example}{\em 
We consider $R=\mathbb F_3[x]/\langle x^{11}-1 \rangle$, and
\[x^{11}-1=(x+2)(x^5 + 2x^3 + x^2 + 2x + 2)(x^5 + x^4 + 2x^3 + x^2 + 2)\in \mathbb F_3[x].\]

We take 

\[g_1(x)=(x+2)(x^5 + x^4 + 2x^3 + x^2 + 2),~~\mbox{and}~~g_2(x)=x^5 + x^4 + 2x^3 + x^2 + 2.\] Then 
\[h_1(x)=(x^5 + 2x^3 + x^2 + 2x + 2),~~\mbox{and}~~h_2(x)=(x+2)(x^5 + 2x^3 + x^2 + 2x + 2).\] 

\noindent We also consider
\[r_1(x)=2x^8 + 2x^7 + 2x^6 + 2x^5 + 2x^4 + 2x^3 + 2x^2 + 2x + 2,~~\mbox{and}\]\[r_2(x) = 2x^7 + 2x^6 + 2x^5 + x^4 + x^3 + 2x^2 + x\in R.\]

\noindent Then  $C=(r_1(x)g_1(x), r_2(x)g_2(x))$ is a one-generator QC code of length $22$ over $\mathbb F_3$. We can check that $\gcd(r_i(x), h_i(x))= 1$  for  $i = 1, 2$. Then $\dim(C)=\deg(h(x))=6$, where \[h(x)=\mathrm{lcm}(h_1(x),h_2(x))=x^6 + 2x^5 + 2x^4 + 2x^3 + x^2 + 1.\]

As per Theorem \ref{t1}, we take $a(x)=r_1(x)g_1(x)\in R,$ and $b(x)=r_2(x)g_2(x)\in R$. Then $A$ is generated by $a(x)$, $B$ is generated by $b(x)$, $A^t$ is generated by $\overline{a}(x)$ and $B^t$ is generated by $\overline{b}(x)$, where 
\[a(x)=x^9 + x^5 + x^4 + x^3 + x + 1,\]
\[\overline{a}(x)=x^{10} + x^8 + x^7 + x^6 + x^2 + 1,\]
\[b(x)=2x^{10} + 2x^8 + 2x^7 + x^6 + x^5 + x^2 + x + 1,\]
\[\overline{b}(x)=x^{10} + x^9 + x^6 + x^5 + 2x^4 + 2x^3 + 2x + 1.\]

\noindent It is easy to check that $AB^t = BA^t$, also \(a(x)\overline{b}(x) - b(x)\overline{a}(x) \equiv 0 \mod (x^n - 1)\). Thus, \(C \subseteq C^{\perp_s}\).
\hfill$\square$
}\end{example}
\vskip 10pt
\begin{remark}
{\em 
A useful property of the transpose polynomial is the following. If $C$ is an $[n, \dim(C)]$ code with circulant generator matrix $G \in \mathbb{F}_q^{n \times n}$ generated by $g(x)$, then the code $C^t$ with generator matrix $G^t$ is also an $[n, \dim(C)]$ code, whose generator is $\overline{g}(x)$. This holds because $\dim(C) = \operatorname{rank}(G) = \operatorname{rank}(G^t) = \dim(C^t)$. Note that the circulant generator matrix need not have full rank.
\hfill$\square$
}\end{remark}

In the following result, we present a theorem that allows us to determine the dimension of a one-generator QC code without the $\gcd$ conditions as in Theorem \ref{gcd1}. A result for an index \( l \) QC code can be found in \cite{La}.

\begin{theorem}
Let \(C\) be a one-generator QC code of length $2n$ and index $2$ generated by \((a(x), b(x))\), where $a(x),b(x)\in R$. Then \(\dim(C) = n - \deg(f(x))\), where \(f(x)=\gcd\big(a(x),b(x),x^n-1\big)\).
\end{theorem}

\begin{proof}
Let $f(x) = \gcd(a(x), b(x), x^n - 1)$.  The code $C$ consists of all codewords whose 
polynomial representations are of the form:
\[
C = \{a(x)p(x) + b(x)q(x) \mid p(x), q(x) \in R\},
\]
so as an ideal of $R$, we have $C = (a(x), b(x))$. We claim that 
$(a(x), b(x)) = (f(x))$ in $R$.

\vskip 5pt
 Since 
$f(x) = \gcd(a(x), b(x), x^n - 1)$, by Bezout's identity there exist polynomials 
$s(x), t(x), w(x) \in \mathbb{F}_q[x]$ such that
\[
f(x) = s(x)a(x) + t(x)b(x) + w(x)(x^n - 1).
\]
Reducing modulo $x^n - 1$, we obtain
\[
f(x) \equiv s(x)a(x) + t(x)b(x) \mod (x^n - 1),
\]
which means $f(x) \in (a(x), b(x))$ in $R$. Hence $(f(x)) \subseteq (a(x), b(x))$.

\vskip 5pt
On the other hand, as  $f(x)$ divides both 
$a(x)$ and $b(x)$ in $\mathbb{F}_q[x]$, we can write $a(x) = f(x)\alpha(x)$ and 
$b(x) = f(x)\beta(x)$ for some $\alpha(x), \beta(x) \in \mathbb{F}_q[x]$. Then 
for any $d(x) \in (a(x), b(x))$, there exist $p(x), q(x) \in R$ such that
\begin{align*}
d(x) &= a(x)p(x) + b(x)q(x) \\
     &= f(x)\alpha(x)p(x) + f(x)\beta(x)q(x) \\
     &= f(x)\bigl(\alpha(x)p(x) + \beta(x)q(x)\bigr).
\end{align*}
Thus $d(x) \in (f(x))$, and therefore $(a(x), b(x)) \subseteq (f(x))$.

\vskip 5pt
Combining both inclusions, we conclude that $(a(x), b(x)) = (f(x))$ in $R$, 
and hence $C = (f(x))$ as an ideal of $R$. Since $f(x)$ divides $x^n - 1$, 
the ideal $(f(x))$ in $R$ is a cyclic code of length $n$ with generator polynomial 
$f(x)$. The dimension of a cyclic code generated by a divisor $f(x)$ of $x^n - 1$ 
is $n - \deg(f(x))$. Therefore,
\[
\dim(C) = n - \deg(f(x)). \qedhere
\]
\end{proof}

\begin{remark}
{\em There are no nontrivial symplectic dual-containing one-generator quasi-cyclic codes. Indeed, the dimension of a one-generator QC code of length $2n$ and index $2$ equals $n - \deg(f(x))$, and consequently the dimension of its symplectic dual is $n + \deg(f(x))$. For the dual-containing condition $C^{\perp_s} \subseteq C$ to hold, we need $\dim(C^{\perp_s}) \le \dim(C)$, i.e., $n + \deg(f(x)) \le n - \deg(f(x))$, which forces $\deg(f(x)) \le 0$, meaning $f(x)$ is a nonzero constant and $C$ is the trivial code $\mathbb{F}_q^{2n}$.\hfill$\square$
}\end{remark}

\section{Two-generators QC codes}
In this section, we present two-generator QC codes, along with the necessary and sufficient conditions for self-orthogonality and the dual-containing property.

In the previous section, we studied one-generator QC codes of the form $(a_1(x), b_1(x))$, where $a_1(x)=r_1(x)g_1(x)$ and $b_1(x)=r_2(x)g_2(x)$, with $g_i(x) \mid x^n - 1$ and $r_i(x) \in R$ for $i = 1, 2$. Similarly, we now introduce two-generator QC codes, where the generators are of the form $(a_1(x), b_1(x))$ and $(a_2(x), b_2(x))$. These generators are defined as follows:
\begin{align}\label{2g1}
    a_1(x)=t_1(x)g_1(x),~~~ b_1(x)=t_2(x)g_2(x),~~~ a_2(x)=t_3(x)g_3(x),~~~ b_2(x) = t_4(x)g_4(x),
\end{align}
where \(a_i(x), b_i(x) \in R\), $g_j(x) \mid x^n - 1$ and $t_j(x) \in R$ for $j = 1, 2, 3, 4$.
 
\vskip 5pt
In this two-generator setting, dealing with four divisors $g_i(x)$ of $x^n-1$ and four auxiliary polynomials $t_i(x) \in R$ can be quite involved. 
Therefore, some special forms have been considered for study. For example, in \cite{Gal}, \cite{Q2}, and \cite{Q4} consecutively, the generators are considered as follows:

\[
 \begin{pmatrix}
f(x) & h(x)f(x)\\
0 & g(x)
\end{pmatrix},
~~\begin{pmatrix}
g_1(x) & g_1(x)\\
g_2(x) & u(x)g_2(x)
\end{pmatrix},~~\mbox{and}~~
\begin{pmatrix}
v(x)g_1(x) & g_1(x)\\
g_2(x) & v(x)g_2(x)
\end{pmatrix}.
\]

\vskip 5pt
In this work, we aim to present a self-orthogonality and dual-containing condition that applies to any two-generator QC codes and can be viewed as a continuation of the one-generator case study. To achieve this, we consider the generator matrix corresponding to the generator \((a_1(x), b_1(x))\) as \(G_1 = (A_1 \mid B_1)\) and for \((a_2(x), b_2(x))\) as \(G_2 = (A_2 \mid B_2)\), where \(A_i\) are circulant matrices generated by the polynomial \(a_i(x)\) for \(i=1,2\), and \(B_i\) are circulant matrices generated by the polynomial \(b_i(x)\) for \(i=1,2\). A generator matrix of the two-generator QC code is then constructed as follows:

\begin{align}\label{2gm}
    G =\begin{pmatrix}
G_1\\
G_2
\end{pmatrix}= \begin{pmatrix}
A_1 & B_1\\
A_2 & B_2
\end{pmatrix}.
\end{align}

Next we give the dimension formula for a  two-generator QC code of length \(2n\) and index \(2\) over \(\mathbb{F}_q\).

\begin{theorem}\label{dim2}
Let \( C \) be a two-generator QC code  of length \(2n\) and index \(2\) over \(\mathbb{F}_q\), with generator matrix \( G \) of the form 
$(\ref{2gm})$ given by

\[ G =\begin{pmatrix}
G_1\\
G_2
\end{pmatrix}.\]
Then, $\dim(C) = \rank(G) = \rank(G_1) + \rank(G_2) - \dim(\text{row space}(G_1) \cap \text{row space}(G_2)).$
\end{theorem}

\begin{remark}{\em 
The result of Theorem \ref{dim2} can also be expressed as 
\[\dim(C) = \dim(C_1) + \dim(C_2) - \dim(C_1 \cap C_2),\]
where we can think \( C_i \) as  one-generator QC codes generated by \( G_i \), for \( i = 1, 2 \). We observed that 
\[\dim(C_i) = \deg(f_i(x)),~\mbox{where}~f_i(x)=\gcd\big(a_i(x), b_i(x), x^n - 1\big),~\mbox{for} ~i=1,2.\]

So far, we have been unable to establish a polynomial degree formula for \(\dim(C_1 \cap C_2)\) in terms of the polynomials considered here. Addressing this issue likely demands further investigation and a more detailed exploration of the polynomials, which we plan to undertake in a future project concentrating on the explicit dual construction of two-generator QC codes.
\hfill$\square$
}\end{remark}

\subsection{Self-orthogonal QC codes}
\begin{theorem}\label{gproof}
Let \(C\) be a two-generator QC code of length \(2n\) and index \(2\) generated by \((a_1(x), b_1(x))\) and \((a_2(x), b_2(x))\), where \(a_i(x), b_i(x) \in R\) and are of the form $(\ref{2g1})$ for \(i = 1, 2\). A generator matrix of this QC code \(C\) is of the form $(\ref{2gm})$. Then \(C \subseteq C^{\perp_s}\) if and only if the following conditions hold:
\[
A_1 B_1^t = B_1 A_1^t, \quad A_2 B_2^t = B_2 A_2^t, \quad \text{and} \quad A_1 B_2^t = B_1 A_2^t,
\]
where \(A_i^t\) and \(B_i^t\) denote the transposes of \(A_i\) and \(B_i\), respectively, for $i=1,2$.
\end{theorem}

\begin{proof}
    By Theorem \ref{mt}, $C$ is a symplectic self-orthogonal code, if and only if  $G\Omega G^t $ is a zero matrix. Here $C$ is a two-generator  QC code of length $2n$ over $\mathbb F_q$, whose generator matrix is  $G$ is of the form $(2)$. Then

\begin{align*}
G \Omega G^t &= \begin{pmatrix}
A_1 & B_1 \\
A_2 & B_2
\end{pmatrix}
\begin{pmatrix}
O_n & I_n \\
-I_n & O_n
\end{pmatrix}
\begin{pmatrix}
A_1^t & A_2^t \\
B_1^t & B_2^t
\end{pmatrix} \\
&= \begin{pmatrix}
-B_1 & A_1 \\
-B_2 & A_2
\end{pmatrix}
\begin{pmatrix}
A_1^t & A_2^t \\
B_1^t & B_2^t
\end{pmatrix} \\
&= \begin{pmatrix}
-B_1A_1^t + A_1B_1^t & -B_1A_2^t + A_1B_2^t\\
-B_2A_1^t + A_2B_1^t & -B_2A_2^t + A_2B_2^t
\end{pmatrix} .
\end{align*}
Therefore,
\[G \Omega G^t = \begin{pmatrix}
-B_1A_1^t + A_1B_1^t & -B_1A_2^t + A_1B_2^t\\
-B_2A_1^t + A_2B_1^t & -B_2A_2^t + A_2B_2^t
\end{pmatrix} = \begin{pmatrix}
O_n & O_n\\
O_n & O_n
\end{pmatrix}.\]

\noindent
Comparing both sides, we obtain the result.
\end{proof}

\begin{remark}{\em
By Theorem~\ref{t1}, the conditions $A_1 B_1^t = B_1 A_1^t$ and $A_2 B_2^t = B_2 A_2^t$ are precisely the symplectic self-orthogonality conditions for the constituent one-generator QC codes $(a_1(x), b_1(x))$ and $(a_2(x), b_2(x))$, respectively. The additional condition $A_1 B_2^t = B_1 A_2^t$ enforces a cross-orthogonality between the two constituent codes. In polynomial terms, by Theorem~\ref{prf}, this is equivalent to $a_1(x)\overline{b}_2(x) - b_1(x)\overline{a}_2(x) \equiv 0 \pmod{x^n-1}$. Thus, the symplectic self-orthogonality of a two-generator QC code reduces to three independent polynomial congruences, each of which can be checked efficiently via polynomial arithmetic in $R$.
\hfill$\square$}\end{remark}

Similar to the one-generator case, we also present an alternative criterion for the symplectic self-orthogonality condition in terms of the generator polynomials for the two-generator QC codes.

\begin{theorem}
Let $C$ be a   two-generator  QC code  of length $2n$ and index $2$ generated by $(a_1(x),b_1(x))$ and $(a_2(x),b_2(x))$,
 where \(a_i(x), b_i(x) \in R\) and are of the form $(\ref{2g1})$ for \(i = 1, 2\). Then $C\subseteq C^{\perp_s}$ if and only if $a_1(x)\overline{b}_1(x)- b_1(x)\overline{a}_1(x) \equiv 0 \mod (x^n - 1)$, $a_2(x)\overline{b}_2(x)-b_2(x)\overline{a}_2(x) \equiv 0 \mod (x^n - 1)$ and $a_1(x)\overline{b}_2(x)-b_1(x)\overline{a}_2(x) \equiv 0 \mod (x^n - 1)$. 
\end{theorem}

\begin{proof}
   This proof follows a similar line of arguments as the proof of Theorem \ref{prf}.
\end{proof}
\subsection{Dual-containing QC codes}
We have examined the symplectic self-orthogonality condition \( C \subseteq C^{\perp_s} \) for two-generator QC codes. Similarly, we can derive a necessary and sufficient condition for the symplectic dual-containing property \( C^{\perp_s} \subseteq C \). Before proceeding, we need the following result.

\begin{theorem}\label{mt2}
Let \( C \) be a linear code of length \( 2n \) over \( \mathbb{F}_q \) with a parity-check matrix \( H \). Suppose \( H \) is an \( m \times 2n \) matrix. Then \( C \) is a symplectic dual-containing code if and only if \( H \Omega H^t = O_m \), where \( O_m \) is the \( m \times m \) zero matrix and \( H^t \) denotes the transpose of \( H \).
\end{theorem}

\begin{proof}
    Let us assume \( C \) is a symplectic dual-containing code, which means \( C^{\perp_s} \subseteq C \). This gives us:
    \begin{align*}
        C^{\perp_s} \subseteq C
       & \Longleftrightarrow \forall x \in C^{\perp_s}, \quad x \in C\\
       & \Longleftrightarrow \forall x \in C^{\perp_s}, \quad H\Omega x^t = O_{m \times 1}\\
       & \Longleftrightarrow \text{for all rows } r \text{ of } H, \quad H\Omega r^t = O_{m \times 1}\\
       & \Longleftrightarrow H\Omega H^t = O_{m}.
    \end{align*}
\end{proof}

To determine the dual-containing property of two-generator QC codes, we need to construct a parity-check matrix. Our objective is to start with a generator matrix \( G \) of a two-generator QC code of length \( 2n \) and index \( 2 \) in the form \((\ref{2gm})\). We consider circulant matrices \( P_i \) generated by the polynomial \( p_i(x) \) for \( i = 1, 2 \), and circulant matrices \( Q_i \) generated by the polynomial \( q_i(x) \) for \( i = 1, 2 \) to form a parity-check matrix \( H \) of the form:
\begin{align} \label{2pm}
H = \begin{pmatrix}
P_1 & Q_1 \\
P_2 & Q_2
\end{pmatrix},
\end{align}

\noindent such that \( G\Omega H^T = O_{2n} \), where \( O_{2n} \) denotes the \( 2n \times 2n \) zero matrix.
\vskip 5pt
By solving \( G\Omega H^T = O_{2n} \), we derive the following equations:
\begin{align*}
A_1 \cdot Q_1^T &= B_1 \cdot P_1^T,\\ 
A_1 \cdot Q_2^T &= B_1 \cdot P_2^T, \\
A_2 \cdot Q_1^T &= B_2 \cdot P_1^T,\\
A_2 \cdot Q_2^T &= B_2 \cdot P_2^T.
\end{align*}

The generator matrix \( G \) in the form \((\ref{2gm})\) and the parity-check matrix \( H \) in the form \((\ref{2pm})\) may not always have full rank. Consequently, \( H \) does not always generate the dual QC code of \( C \). The condition \( G\Omega H^T = O_{2n} \) indicates that if a two-generator QC code \( C \) is generated by the matrix \( G \) in the form \((\ref{2gm})\), and another two-generator QC code \( C' \) is generated by the matrix \( H \) in the form \((\ref{2pm})\), then all codewords of \( C \) are orthogonal to those of \( C' \). However, \( C' \) is not always equal to \( C^{\perp_s} \), the symplectic dual of \( C \). If the matrix \( H \) satisfies \(\text{rank}(G) + \text{rank}(H) = 2n\), i.e., \( \dim(C) + \dim(C') =2n\), we can assert that \( H \) is the parity-check matrix of \( C \) that generates \( C^{\perp_s} \).
\vskip 5pt

Suppose  $g(x)$ is  a monic  polynomial such that $g(x)h(x)=x^n -1$. Then the polynomial $g^\perp(x)$  is defined as 
\[g^\perp(x)=x^{\deg(h)}h(x^{-1})/h(0).\]

\begin{example}\label{e1}{\em
We consider $R=\mathbb{F}_3[x]/\langle x^{15}-1 \rangle$, and
\[
x^{15}-1 = (x+2)^3(x^4 + x^3 + x^2 + x + 1)^3 \in \mathbb{F}_3[x].
\]

We take 
\[g_1(x) = (x+2)(x^4 + x^3 + x^2 + x + 1), ~~~~g_2(x) = (x+2)^3(x^4 + x^3 + x^2 + x + 1),\]
\[g_3(x) = (x^4 + x^3 + x^2 + x + 1)^2,~~\mbox{and}~~
g_4(x) = (x^4 + x^3 + x^2 + x + 1)\]
such that $g_i(x) \mid x^{15}-1$ for $i=1,2,3,4$. We also take $t_i(x) \in R$ for $i=1,2,3,4$ such that
\begin{align*}
t_1(x) &= 2x^{12} + 2x^{11} + 2x^{10} + 2x^9 + 2x^8 + 2x^7 + 2x^6 + 2x^5 + 2x^4 + 2x^3 + 2x^2 + 2x + 2, \\
t_2(x) &= 2x^{11} + 2x^{10} + 2x^9 + 2x^8 + 2x^7 + 2x^6 + 2x^5 + 2x^4 + 2x^3 + 2x^2 + 2x + 2, \\
t_3(x) &= 2x^9 + 2x^8 + 2x^7 + 2x^6 + 2x^5 + 2x^4 + 2x^3 + 2x^2 + 2x + 2, \\
t_4(x) &= 2x^7 + 2x^6 + 2x^5 + 2x^4 + 2x^3 + 2x^2 + 2x + 2.
\end{align*}

\noindent Then \(C\) is a two-generator QC code of length \(30\) and index \(2\) generated by \((a_1(x), b_1(x))\) and \((a_2(x), b_2(x))\), where 
$a_i(x) \equiv t_i(x)g_i(x) \mod (x^{15}-1)$ for $i=1,2$ and $b_j(x) \equiv  t_j(x)g_j(x) \mod (x^{15}-1)$ for $j=3,4$. Then the generator matrix $G$ is of the form $(\ref{2gm})$. 

\vskip 3pt

We consider $p_i(x) \equiv  g^\perp_i(x)\overline{t}_i(x) \mod (x^{15}-1)$ for $i=1,2$ and $q_j(x) \equiv  g^\perp_j(x)\overline{t}_j(x) \mod (x^{15}-1)$ for $j=3,4$. Then the parity-check matrix $H$ is of the form \((\ref{2pm})\). We can check that  \( G \Omega H^T = O_{30} \) and \(\text{rank}(G) + \text{rank}(H) = 30\). Thus $H$ generates the symplectic dual of the QC code $C$.
}\hfill$\square$
\end{example}

\vskip 5pt
Assuming that a parity-check matrix of the form~(\ref{2pm}) generates $C^{\perp_s}$, we derive the following necessary and sufficient condition for a two-generator QC code to be symplectic dual-containing.

\begin{theorem}\label{1}
Let \(C\) be a two-generator QC code of length \(2n\) and index \(2\) over $\mathbb F_q$. A parity-check matrix of this QC code \(C\) is of the form $(\ref{2pm})$. Then \(C^{\perp_s} \subseteq C\) if and only if the following conditions hold:
\[
P_1 Q_1^t = Q_1 P_1^t, \quad P_2 Q_2^t = Q_2 P_2^t, \quad \text{and} \quad P_1 Q_2^t = Q_1 P_2^t,
\]
where \(P_i^t\) and \(Q_i^t\) denote the transposes of \(P_i\) and \(Q_i\), respectively, for \(i=1,2\).
\end{theorem}

\begin{proof}
    The proof follows a similar approach to the proof of Theorem \ref{gproof}.
\end{proof}

A necessary and sufficient condition for two-generator QC codes of length \(2n\) and index \(2\) over \(\mathbb{F}_q\) that contain their duals can be expressed in terms of polynomials as follows.

\begin{theorem}\label{2}
Let \(C\) be a two-generator QC code of length \(2n\) and index \(2\) over $\mathbb F_q$. 
A parity-check matrix of this QC code \(C\) is of the form $(\ref{2pm})$.
Then \(C^{\perp_s} \subseteq C\) if and only if $
p_1(x)\overline{q}_1(x) - q_1(x)\overline{p}_1(x) \equiv  0 \mod (x^n - 1), p_2(x)\overline{q}_2(x) - q_2(x)\overline{p}_2(x) \equiv  0 \mod (x^n - 1),$ and $p_1(x)\overline{q}_2(x) - p_2(x)\overline{q}_1(x) \equiv  0 \mod (x^n - 1),$ where \(\overline{p}(x)\) denotes the transpose polynomial of \(p(x)\).
\end{theorem}
\begin{proof}
   This proof follows a similar line of arguments as the proof of Theorem \ref{prf}.
\end{proof}

\begin{example}{\em 
Continuing from Example \ref{e1}, we can demonstrate that the two-generator QC code $C$ described in Example \ref{e1} meets both the necessary and sufficient conditions for the dual-containing property as stated in Theorem \ref{1} and Theorem \ref{2}. Consequently, this code is a dual-containing QC code of length $2n$ and index $2$ over $\mathbb{F}_3$.\hfill$\square$
}\end{example}

\begin{remark}{\em 
The duals of one-generator QC codes have been addressed in \cite{QC, iitr}. The structure of duals of two-generator quasi-cyclic codes is considerably more complex, primarily because the generator matrix involves eight polynomials. In this work, we identify symplectic self-orthogonal and symplectic dual-containing codes without explicitly computing the generators of the dual code. Minimum distance bounds for specific two-generator QC codes have been discussed in \cite{Gal, Q2}; establishing such bounds for \emph{general} two-generator QC codes remains an open problem. 
\hfill$\square$}\end{remark}

\section{QECCs from QC codes}
Most quantum codes studied in the literature are based on the well-known CSS construction \cite{CRSS98}. Quantum codes over $\mathbb{F}_q$ (where $q$ is a prime power) have also been constructed using Hermitian and symplectic inner products, as studied in \cite{A, Ket}. We recall the quantum code construction result from \cite{Ket} based on symplectic self-orthogonality, and state its analogue for the dual-containing property.
\vskip 3pt
To construct a quantum code from a linear code $C$ via the symplectic framework, one must ensure that either $C \subseteq C^{\perp_s}$ or $C^{\perp_s} \subseteq C$. The primary goal of this paper is to establish necessary and sufficient conditions that allow efficient construction of linear codes with the symplectic self-orthogonal or symplectic dual-containing property. Based on these two properties, we derive the corresponding results that we  use to construct quantum codes from our study.

\begin{theorem}
(\cite{0}) \label{Q}
Let \( C \) be a linear code of length \( 2n \) over \( \mathbb{F}_q \) with parameters \([2n, k]\). If \( C \subseteq C^{\perp_s} \), then there exists a quantum error-correcting code \( Q \) with parameters \( [[n, n-k, d_s]] \) over \( \mathbb{F}_q \), where \( d_s = \min\{w_s(c) \mid c \in (C^{\perp_s} \setminus C)\} \).
\end{theorem}

We can also state the above result in terms of the dual-containing property.

\begin{theorem}\label{dc}
Let \( C \) be a linear code of length \( 2n \) over \( \mathbb{F}_q \) with parameters \([2n, k]\). If \( C^{\perp_s} \subseteq C \), then there exists a quantum error-correcting code \( Q \) with parameters \( [[n, k-n, d_s]] \) over \( \mathbb{F}_q \), where \( d_s = \min\{w_s(c) \mid c \in (C \setminus C^{\perp_s})\} \).
\end{theorem}

\begin{proof}
Let \( C \) be a linear code of length \( 2n \) over \( \mathbb{F}_q \) with parameters \([2n, k]\), such that \( C^{\perp_s} \subseteq C \). Consider \( D = C^{\perp_s} \), which is a linear code of length \( 2n \) over \( \mathbb{F}_q \) with parameters \([2n, 2n-k]\). Since \( D = C^{\perp_s} \), it follows that \( D^{\perp_s} = C \). Therefore, \( C^{\perp_s} \subseteq C \) implies \( D \subseteq D^{\perp_s} \).
\vskip 3pt
By Theorem \ref{Q}, there exists a quantum error-correcting code \( Q \) with parameters \( [[n, n-(2n-k), d_s]] = [[n, k-n, d_s]] \), where \( d_s = \min \{ w_s(c) \mid c \in C \setminus C^{\perp_s} \} \).
\end{proof}

For ease of computation, we primarily consider one-generator quasi-cyclic codes \( C \) of the form \((r_1(x)g(x), r_2(x)g(x))\), where \( g(x) \mid x^n - 1 \) and \( r_1(x), r_2(x) \in R \). We observe that a quantum code constructed from a symplectic self-orthogonal QC code \( C \) of this form has a degree given by \(n-k = n-(n-\deg(g(x)))=\deg(g(x))\). The advantage of this choice is that it allows us to fix the degree of the quantum code to match the dimension parameter of the code we seek to improve. All computations are done using MAGMA \cite{Mag}. 

\begin{example}{\em
We consider \( q = 5 \) and \( n = 11 \). Then \( R = \mathbb{F}_5[x]/\langle x^{11} - 1 \rangle \). We take two polynomials \( r_1(x), r_2(x) \in R \), where
\[ r_1(x) = 4x^8 + 4x^7 + 4x^6 + 4x^5 + 4x^4 + 4x^3 + 4x^2 + 4x + 4, \]
\[ r_2(x) = 4x^6 + 2x^5 + 4x^4 + 2x^3 + 4x^2, ~~g(x) = 1. \]
Next, we consider two circulant matrices of size 11, \( A \) and \( B \), generated by \( r_1(x) \) and \( r_2(x) \) over \( \mathbb{F}_5 \). Then \( C \) is a QC code of length $22$ and index $2$ whose generator matrix is \( G = (A \mid B) \), where \(\mid\) represents the horizontal concatenation of the two circulant matrices \( A \) and \( B \). We note that \( AB^t = BA^t \), which implies \( C \) is a symplectic self-orthogonal code with parameters \([22, 11, 8]\) over \( \mathbb{F}_5 \). Therefore, by Theorem \(\ref{Q}\), we obtain a QECC with parameters \([[11, 0, 6]]\), which are record-breaking parameters. The previous record was \([[11, 0, 5]]\). This  code has been already updated to the  quantum code table \cite{gra}.
\hfill$\square$}\end{example}

\begin{example}{\em
We consider \( q = 3 \) and \( n = 13 \). Then \( R = \mathbb{F}_3[x]/\langle x^{13} - 1 \rangle \). We take $g(x)\mid x^{13}-1$ and \(  r_1(x), r_2(x) \in R \), where
\[ g(x) = 2x^6 + x^5 + 2x^4 + x^3 + x^2 + x + 2, \]
\[ r_1(x) = 2x^6 + 2x^5 + 2x^4 + 2x^3 + 2x^2 + 2x + 1, \]
\[ r_2(x) = 2x^6 + 2x^5 + 2x^4 + x^3 + 2x^2 + 2x + 2. \]
Next, we consider two circulant matrices of size $13$, \( A \) and \( B \), generated by \( g(x)r_1(x) \) and \( g(x)r_2(x) \) over \( \mathbb{F}_3 \). Then \( C \) is a QC code of length $26$ with index $2$ whose generator matrix is \( G = (A \mid B) \), where \(\mid\) represents the horizontal concatenation of the two circulant matrices \( A \) and \( B \). We note that \( AB^t = BA^t \), which implies \( C \) is a symplectic self-orthogonal code with parameters \([26, 7, 12]\) over \( \mathbb{F}_3 \). By Theorem \(\ref{Q}\), we obtain a QECC with parameters \([[13, 6, 4]]\), which are record-breaking parameters. The previous record was \([[13, 6, 3]]\). This  code has been already updated to the  quantum code table \cite{gra}.
\hfill$\square$}\end{example}

\begin{example}{\em
We consider \( q = 3 \) and \( n = 23 \). Then \( R = \mathbb{F}_3[x]/\langle x^{23} - 1 \rangle \). We take $g(x)\mid x^{23}-1$ and \(  r_1(x), r_2(x) \in R \), where
\[ g(x) = x^{12} + 2x^{11} + 2x^9 + x^8 + 2x^7 + x^6 + x^5 + x^3 + 1, \]
\[ r_1(x) = 2x^{11} + 2x^{10} + 2x^9 + 2x^8 + 2x^7 + 2x^6 + 2x^5 + 2x^4 + 2x^3 + 2x^2 + 2x + 2, \]
\[ r_2(x) = 2x^{10} + 2x^9 + 2x^8 + 2x^7 + 2x^6 + 2x^5 + 2x^4 + x^3 + 2x^2 + 1. \]
Next, we consider two circulant matrices of size 23, \( A \) and \( B \), generated by \( g(x)r_1(x) \) and \( g(x)r_2(x) \) over \( \mathbb{F}_3 \). Then \( C \) is a QC code of length $46$ with index $2$ whose generator matrix is \( G = (A \mid B) \), where \(\mid\) represents the horizontal concatenation of the two circulant matrices \( A \) and \( B \). We note that \( AB^t = BA^t \), which implies \( C \) is a symplectic self-orthogonal code with parameters \([46, 11, 21]\) over \( \mathbb{F}_3 \). By Theorem \(\ref{Q}\), we obtain a QECC with parameters \([[23, 12, 5 ]]\), which are record-breaking parameters. The previous record was \([[23, 12, 4 ]]\). This  code has been already updated to the  quantum code table \cite{gra}.
\hfill$\square$}\end{example}

\begin{example}{\em
We consider \( q = 3 \) and \( n = 16 \). Then \( R = \mathbb{F}_3[x]/\langle x^{16} - 1 \rangle \). We take $g(x)\mid x^{16}-1$ and \(  r_1(x), r_2(x) \in R \), where
\[ g(x) = 2x^6 + x^4 + 1, \]
\[ r_1(x) = 2x^9 + 2x^8 + 2x^7 + 2x^6 + 2x^5 + 2x^4 + 2x^3 + 2x^2 + 2x + 1, \]
\[ r_2(x) = 2x^9 + 2x^8 + x^7 + 2x^6 + x^5 + x^4 + 2x^3 + x. \]
Next, we consider two circulant matrices of size $16$, \( A \) and \( B \), generated by \( g(x)r_1(x) \) and \( g(x)r_2(x) \) over \( \mathbb{F}_3 \). Then \( C \) is a QC  code of length $32$ with index $2$ whose generator matrix is \( G = [A \mid B] \), where \(\mid\) represents the horizontal concatenation of the two circulant matrices \( A \) and \( B \). We note that \( AB^t = BA^t \), which implies \( C \) is a symplectic self-orthogonal code with parameters \([32, 10, 12]\) over \( \mathbb{F}_3 \). By Theorem \(\ref{Q}\), we obtain a QECC with parameters \([[16, 6, 5]]\), which are record-breaking parameters.
The previous record was \([[16, 6, 4 ]]\). This  code has been already updated to the  quantum code table \cite{gra}.
\hfill$\square$}\end{example}

\begin{example}{\em
By \cite[Theorem 6]{cal1996}, if  a quantum code with parameters   $ [[n,k,d]]$ exists then a quantum code with parameters $ [[n+1,k,d]]$ also exists, when $k > 0$. Therefore, from the above-constructed quantum code parameters  $ [[ 16, 6, 5 ]]$, we get a quantum code with parameters  $ [[ 17, 6, 5 ]]$ which is also new and breaks the previous record which is $ [[ 17, 6, 4 ]]$. This newly constructed code is  in the online quantum code table \cite{gra}.
\hfill$\square$}\end{example}

\section{Conclusion and Future work} 

In this work, we have studied one-generator and two-generator quasi-cyclic (QC) codes over \(\mathbb{F}_q\), where \(q\) is a prime power. We presented necessary and sufficient conditions for the symplectic self-orthogonality of one-generator QC codes in both matrix and polynomial forms, and used these to construct new quantum codes with record-breaking parameters. Extending this study to two-generator QC codes, we presented necessary and sufficient conditions for both symplectic self-orthogonality and the symplectic dual-containing property. For each factor \(g(x)\) of \(x^n - 1\), we choose two polynomials \(r_1(x)\) and \(r_2(x)\) to construct a quantum code from the one-generator QC codes. Unlike the commutative polynomial ring $\mathbb{F}_q[x]$, skew polynomial rings are not unique factorization domains; as a result, a given element can admit multiple inequivalent factorizations. This multiplicity increases the number of potential generator polynomials and hence the opportunity to construct new codes. This multiplicity increases the potential to find more factors and, consequently, more possibilities to construct codes. It would be interesting to study one-generator skew quasi-cyclic codes. Extend the symplectic self-orthogonality and dual-containing conditions to one-generator and two-generator skew quasi-cyclic codes over $\mathbb{F}_q$, exploiting the non-unique factorization in skew polynomial rings to produce new record-breaking quantum codes.

\section*{Acknowledgement}
Part of this work was carried out while T. Bag was with Inria, ENS de Lyon and supported by the European Research Council (ERC Grant AlgoQIP, Agreement No. 851716) and by the Agence Nationale de la Recherche (ANR) under the France 2030 program, grant ANR-22-PETQ-0006. T. Bag is grateful to Prof. Markus Grassl for numerous discussions on quantum codes. D. Panario was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC), grant number RPGIN-2018-05328. We sincerely thank the reviewer and editor for their valuable and constructive feedback, which greatly helped improve the clarity and overall quality of this work.

\section*{Conflict of interest: } The authors declare that there are no conflicts of interest.

\end{document}